I need help solving each system by addition, and determine whether the equations are independent, dependent, or inconsistent. Here is the problem.
x- y=3
-6x+6y=17.

Answer by stanbon(75887)

determine whether the equations are independent, dependent, or inconsistent. Here is the problem.
x- y=3
-6x+6y=17.
------
Multiply thru the 1st equation by 6::
6x - 6y = 18
-6x + 6y = 17
------
Add to get:
0 = 35
That is a contradiction::
----
Note: This contradiction happened because I assumed
the x's and the y's in the two equations had the same variable value.
------
The system is inconsistent.
=================